There are 30 employees in total in the company, 5 of them are males and 25 are females. Two employees are selected to participate in a quality assurance survey.

What is the probability of selecting a male employee followed by another male employee?
*Round your answer to 4 decimal places.
Answer #1

P(Male ) = 5 / 30

Since first male is already selected , there are 3 male employees left with 29 total employees.

P(second male) = 4/29

P(male followed by male) = 5/30 * 4/29

= 0.0230